OpenStudy (anonymous):

Given f(x) =x-5/3 , solve for f-1(3). (5 points) Question 2 options: 1) -1 2) 14 3) 18 4) 21

OpenStudy (johnweldon1993):

\[\large f(x) = x - \frac{5}{3}\] right?

OpenStudy (anonymous):

Given f(x) =x-5/3 , solve for f^-1(3). (5 points) Question 2 options: 1) -1 2) 14 3) 18 4) 21

OpenStudy (anonymous):

its to the -1 power I wrote it wrong

OpenStudy (johnweldon1993):

I got what you meant though, but I'm making sure the f(x) equation is correct for now Is it all over 3...or is it as I have written above?

OpenStudy (anonymous):

(x-5)/3 @johnweldon1993

OpenStudy (shamim):

Let f(x)=y

OpenStudy (shamim):

So y=(x-5)/3 3y=x-5 x=3y+5

OpenStudy (shamim):

So f^-1(y)=3y+5

OpenStudy (shamim):

Tell me f^-1(x)=?

OpenStudy (shamim):

Just replace y by x